OpenStudy (anonymous):

The Question is: Write an equation for the line that is parallel to the given line and passes through the given point. Write your equation in slope-intercept form. y = 4x - 3; (-3,12). I don't want the answer, I want to know how to figure it out. Any help is appreciated :)

OpenStudy (anonymous):

tell me for two lines if the slopes are parallel what would their slopes be in relation to each other?

OpenStudy (anonymous):

The reciprocal right?

OpenStudy (anonymous):

no negative reciprocal is for lines perpendicular to each other

OpenStudy (anonymous):

the slope would be the same for lines that are parallel to each other

OpenStudy (anonymous):

first of all from what they give u tell me what is the slope?

OpenStudy (anonymous):

According to the given equation it would be 4/1.

OpenStudy (anonymous):

yes and so now u have a m=4 and the pt given (-3,12) which represents (x,y) using y=mx+b solve for b

OpenStudy (texaschic101):

nice explanation...ngott :)

OpenStudy (anonymous):

The y intercept would be 24 I think.

OpenStudy (anonymous):

assuming ur algebra is right now u have everything u need to write the equation in slope intercept form y=4x+24

OpenStudy (anonymous):

That really helped :). I only have one more problem on the worksheet like this one. I think I can do it though :P

OpenStudy (texaschic101):

y = mx + b (m = 4, x = -3, y = 12) 12 = 4(-3) + b 12 = -12 + b 12 + 12 = b 24 = b yes...your equation is : y = 4x + 24

OpenStudy (texaschic101):

if you were looking for the slope intercept form, you could use y = mx + b to solve it, or you could use y - y1 = m(x - x1) and then put it in y = mx + b form. But if it specifically asks for slope intercept form, would just use y - y1 = m(x - x1)

OpenStudy (texaschic101):

oops...specifically asks for point slope form, you would use y - y1 = m(x - x1)

OpenStudy (anonymous):

@Hardcorefr3k im sorry but @texaschic101 is correct